# Monotone Neural Barrier Certificates

### v1

@@ Abstract
This work presents a neurosymbolic framework for safety verification and control synthesis in high-dimensional monotone dynamical systems, using an algorithm with linear sample complexity—without relying on explicit models or conservative Lipschitz bounds. The approach combines the expressiveness of neural networks with the rigor of symbolic reasoning via barrier certificates, functional analogs of inductive invariants that formally guarantee safety. Prior data-driven methods often treat dynamics as black-box models, relying on dense state-space discretization or Lipschitz overapproximations, leading to exponential sample complexity. In contrast, monotonicity—a pervasive structural property in many real-world systems—provides a symbolic scaffold that simplifies both learning and verification. Exploiting order preservation reduces verification to localized boundary checks, transforming a high-dimensional problem into a tractable, low-dimensional one. Barrier certificates are synthesized using monotone neural networks—architectures with embedded monotonicity constraints—trained via gradient-based optimization guided by barrier conditions. This enables scalable, formally sound verification directly from simulation data, bridging black-box learning and formal guarantees within a unified neurosymbolic framework. Empirical results on three large-scale benchmarks—a freeway traffic model with increasing size, a 50-dimensional urban traffic network, and a 13,000-dimensional power grid—demonstrate the scalability and effectiveness of the approach in real-world, safety-critical systems.
1 Introduction
@@
Discrete-time dynamical systems offer a foundational framework for modeling and controlling a broad spectrum of digitally controlled autonomous systems, including autonomous vehicles, implantable medical devices, biological networks, and industrial process management. As these systems constitute a critical component of our safety infrastructure, high-assurance formal reasoning is essential to ensure their safe operation. A principled and efficient approach to providing such safety guarantees involves synthesizing barrier certificates prajna2006barrier ; ADA-SC-ME-GN-KS-PT:19 —a functional analog of inductive invariants—whose existence serves as a formal proof of safety AP-AR:04 . Although finding such certificates is inherently computationally challenging, recent neuro-symbolic approaches have shown considerable promise through the use of neural barrier certificates chang2019neural ; zhao2020synthesizing ; zhang2024exact ; robey2020learning . While these neural certificates provide substantial flexibility in searching for safety guarantees, we posit that scalability and sample efficiency can be further improved in specialized control applications by identifying and leveraging structural properties of the underlying system. In this work, we focus on a subclass of discrete-time dynamical systems known as monotone systems angeli2003monotone ; HLS:95 , which frequently arise in domains such as queuing networks farina2011positive , population dynamics leenheer2004predator , biological systems DA:13 , and chemical networks EDS:07 . By leveraging their structural characteristics, we synthesize barrier certificates represented as monotone neural networks, which we refer to as monotone neural barrier certificates.
@@
Monotone Systems. Monotone dynamical systems HLS:95 ; angeli2003monotone are characterized by the property that their trajectories preserve a natural partial order induced on their state set. As a paradigmatic example, consider traffic flow dynamics sadraddini2018formal , where increasing the number of vehicles on a road segment does not lead to a decrease in vehicle density downstream under normal flow conditions. That is, if one traffic configuration contains more vehicles than another, the evolution of the system preserves this ordering. Similar examples arise in domains such as biochemical reactions in synthetic biology EDS:07 , frequency regulation in power system, congestion control in transportation networks SC-MA:15 ; GC-EL-KS:15 , population dynamics in ecology, and compartmental models in epidemiology angeli2003monotone . Related concepts have also been studied in other domains, including positive linear systems farina2011positive , well-structured transition systems FINKEL200163 , and well-behaved transition systems Blondin2016WellBT . It is well known that monotone systems exhibit highly ordered transient and asymptotic behaviors HLS:95 , making them particularly well-suited for data-driven methods, as their inherent structure helps mitigate the challenges posed by limited or noisy data. While monotonicity has informed control and stability analysis AR:15 ; SC:19 , its potential for learning-based safety verification remains largely untapped.
@@
Barrier Certificates (A Proof of Safety). In safety verification, the central object of interest is the reachable set, which comprises all states the system can reach from a given set of initial states under a fixed control policy. Safety can be certified by showing that this set does not intersect the set of harmful or unsafe states. However, computing the exact reachable set is generally intractable for nonlinear or high-dimensional systems. A common alternative is to construct an inductive state-invariant, i.e., a set that contains the initial states and is closed under the system’s transitions. Barrier certificates prajna2006barrier offer one such approach by defining a real-valued function over the state space whose level sets form a “barrier” separating safe and unsafe regions. Specifically, the barrier function must: 1) take values less than or equal to a constant over the initial states, 2) and strictly greater than that constant over unsafe states; and 3) decrease along system trajectories outside the unsafe set. A control barrier certificate (CBC) generalizes the idea of barrier certificate by requiring the existence of an action satisfying the conditions for a barrier certificate.
@@
Despite their intuitive appeal and ease of verification, complexity of finding CBCs remains a significant hurdle in their widespread adoption. Traditional methods for synthesizing CBCs rely on optimization-based formulations such as Sum-of-squares (SOS) programming AP-SJ:05 ; kong2013exponential . SOS is a popular approach that is applicable to systems with polynomial dynamics; however, it quickly becomes intractable due to the growth of semidefinite constraints with system dimension and polynomial degree. Linear programming (LP) methods, while more tractable, are typically limited to linear or affine certificates and cannot capture the rich nonlinear boundaries present in many real-world systems SS-XC-EA:13 ; YZ-etal:16 . Recent progress in deep learning has enabled the use of neural networks to be trained as barrier certificates anand2023formally ; lindemann2024learning ; dawson2022safe ; liu2023safe ; zhao2020synthesizing ; Rus2023 ; zhang2024exact . These so-called neural control barrier certificates (NCBCs) represent both the barrier function and controller as neural networks. While expressive, these approaches often require full knowledge of the system dynamics (to verify correctness using constraint satisfaction tools such as Z3 de2008z3 ) or rely on bounding Lipschitz constants (which entails exponential sample complexity), limiting scalability and applicability to black-box systems anand2023formally ; zhang2024exact . Moreover, these approaches treat the system as unstructured, ignoring properties like monotonicity that could drastically reduce the number of required samples and constraints [^1] We refer to Appendix A for a comprehensive review of related works..
@@
Monotone Neural Control-Barrier Certificates. We introduce a machine learning algorithm with linear sample complexity for synthesizing control barrier certificates specifically tailored to (unknown) monotone systems. By exploiting monotonicity, we restrict the satisfaction of safety conditions to critical boundary slices of the state space—namely, the upper boundary of the initial set and the lower boundary of the unsafe set. This reduces the verification burden from the entire state space to a carefully selected subset—without compromising formal guarantees—and enables the use of gradient-based optimization to learn barrier certificates using simulation data alone. Our method avoids the need for explicit models, Lipschitz bounds, or semidefinite relaxations, and yields an algorithm for synthesizing barrier certificates that scales linearly with the system dimension. This makes it uniquely suited for safety verification in large-scale, nonlinear systems where monotonicity arises naturally from physical or structural properties.
Contributions. In summary, this work introduces a learning framework for safety verification and safe control synthesis of unknown monotone dynamical systems using monotone neural barrier certificates. The key contributions of our paper are as follows:
  1. 1.
    Monotone structure of barrier certificates. We investigate the structure of barrier certificates for monotone dynamical systems and establish a theoretical result demonstrating that, under mild conditions on the unsafe set, monotone barrier certificates can effectively characterize the system’s safety. This result shows that restricting to monotone certificates incurs no loss of generality and supports their use to improve sample complexity and computational cost of safety verification.
  2. 2.
    Monotonicity as a catalyst for scalable verification. By leveraging the monotone structure of barrier certificates, we propose a scalable framework that verifies safety using only a finite amount of simulator data—without requiring explicit models or Lipschitz constants—by reducing verification to localized boundary checks.
  3. 3.
    Monotone Neural Barrier Certificates. We propose an algorithm that employs monotone neural networks to learn barrier certificates by leveraging reduced safety conditions on the boundaries of the state set. Unlike existing methods based on sum-of-squares, linear programming, or neural barrier certificates, our algorithm achieves linear sample complexity—rather than exponential—and supports arbitrary partitioning of the state set.
  4. 4.
    Validation on large-scale, safety-critical domains. We evaluate the efficiency of our algorithm on three challenging case studies: (i) a 1,000-dimensional freeway traffic model, (ii) a 50-dimensional urban traffic control network, and (iii) a 13,000-dimensional power grid. These results demonstrate the practicality and scalability of our method for safety-critical real-world applications. Moreover, in the high-dimensional freeway traffic model, we show that our algorithm outperforms state-of-the-art methods for synthesizing barrier certificates.
2 Preliminaries and Mathematical Foundations
@@
We denote the set of real, non-negative real, and positive real numbers by $\mathbb{R}$, $\mathbb{R}_{\geq 0}$, and $\mathbb{R}_{>0}$, respectively. For sets $A$ and $B$, we use $A\setminus B$ for set difference and $A\times B$ for Cartesian product, and write $|A|$ for the cardinality of $A$. Given a vector $v=(v_{1},\ldots,v_{n})^{\top}\in\mathbb{R}^{n}$, its Euclidean norm is $\|v\|=\sqrt{\sum_{i=1}^{n}v_{i}^{2}}$. The rectified linear unit is defined by $\mathsf{ReLU}(x):=\max(x,0)$, and for vectors $x,y\in\mathbb{R}^{n}$, the mean squared error is $\mathsf{MSE}(x,y):=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-y_{i})^{2}$. We denote the $i$th component of a vector $x$ by $x_{i}$, and write $x\leq y$ to indicate element-wise inequality. A set $\mathcal{S}\subseteq\mathbb{R}^{n}$ is lower closed if $x\in\mathcal{S}$ and $y\leq x$ imply $y\in\mathcal{S}$; it is upper closed if $x\in\mathcal{S}$ and $y\geq x$ imply $y\in\mathcal{S}$. A hyper-rectangle is denoted by $\underline{x},\overline{x}]:=\{x\in\mathbb{R}^{n}\mid\underline{x}\leq x\leq\overline{x}\}$. Given a set $A\subseteq^{n}$, $x\in^{n}$ is a maximal (resp. minimal) point of $A$, if there does not exists any $y\in A$ such that $x\leq y$ (resp. $y\leq x$).
2.1 Safety of Dynamical Systems
@@
Definition 1 (Discrete-time dynamical system).
@@
$x_{t+1}=f(x_{t},u_{t}),\qquad\text{for all }t\geq 0.$ (1)
The sequence $\{x_{t}\}_{t=0}^{\infty}$ defined by (1) is called a trajectory of $\mathfrak{S}$ for the input $\{u_{t}\}_{t=0}^{\infty}$. Given a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$, a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$, and an input sequence $\{u_{t}\}_{t=0}^{\infty}$, we say that $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$ for this input sequence if all trajectories starting from initial states in $\mathcal{X}_{0}$ remain outside of $\mathcal{X}_{u}$. That is, for every $x_{0}\in\mathcal{X}_{0}$, the corresponding trajectory $\{x_{t}\}_{t=0}^{\infty}$ for the input sequence $\{u_{t}\}_{t=0}^{\infty}$ starting from $x(0)=x_{0}$, satisfies
@@
In safety verification, inputs are viewed as disturbances, and the goal is to guarantee system safety under all possible input sequences $\{u_{t}\}_{t=0}^{\infty}\in U$. To this end, robust barrier certificates have been widely employed as an effective tool for safety verification (prajna, ).
Definition 2 (Robust barrier certificates).
@@
$B(x)\leq\gamma,\qquad\text{ for all }x\in\mathcal{X}_{0},$ (2)
$B(x)>\eta,\qquad\text{ for all }x\in\mathcal{X}_{u},\text{ and }$ (3)
$B(x)\leq\delta\implies\mathrm{sup}_{u\in U}\{B(f(x,u))\}\leq\delta.$ (4)
The following theorem formulates the safety verification problem as a search for RBCs prajna .
Theorem 1 (Safety verification via robust barrier certificates).
Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ and a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Suppose that there exists a robust barrier certificate satisfying conditions (2)-(4). For any input sequence $\{u_{t}\}_{t=0}^{\infty}\in U$, the system $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$.
Proof. We refer to Appendix B for the proof.
@@
In safe control synthesis, inputs are treated as control policies, aiming to synthesize a controller $\{u_{t}\}_{t=0}^{\infty}$ that ensures system safety. Control barrier certificates have been widely adopted as an effective tool for synthesizing such controllers (ADA-XX-JWG-PT:17, ).
Definition 3 (Control barrier certificates).
Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. A function $B:\mathcal{X}\rightarrow\mathbb{R}$ is called a Control Barrier Certificate (CBC) for $\mathfrak{S}$ with respect to the unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$ if there exists $\gamma,\delta,\eta\in\mathbb{R}_{\geq 0}$ such that $\gamma\leq\delta\leq\eta$ and $B$ satisfy conditions (2), (3), and the following condition
$B(x)\leq\delta\implies\mathrm{inf}_{u\in U}\{B(f(x,u))\}\leq\delta.$ (5)
The next result formulates the safe control synthesis problem as a search for CBCs ADA-XX-JWG-PT:17 .
Theorem 2 (Control synthesis via control barrier certificates).
Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ and a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Suppose that there exists a control barrier certificate satisfying conditions (2), (3), and (5). Then, there exists a map $K:\mathcal{X}\to U$ such that the system $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
Proof. We refer to Appendix C for the proof.
2.2 Monotone Dynamical Systems

### v3

@@ Abstract.
This report presents a neurosymbolic framework for safety verification and control synthesis in high-dimensional monotone dynamical systems without relying on explicit models or conservative Lipschitz bounds. The approach combines the expressiveness of neural networks with the rigor of symbolic reasoning via barrier certificates, functional analogs of inductive invariants that formally guarantee safety. Prior data-driven methods often treat dynamics as black-box models, relying on dense state-space discretization or Lipschitz overapproximations, leading to exponential sample complexity. In contrast, monotonicity—a pervasive structural property in many real-world systems—provides a symbolic scaffold that simplifies both learning and verification. Exploiting order preservation reduces verification to localized boundary checks, transforming a high-dimensional problem into a tractable, low-dimensional one. Barrier certificates are synthesized using monotone neural networks—architectures with embedded monotonicity constraints—trained via gradient-based optimization guided by barrier conditions. This enables scalable, formally sound verification directly from simulation data, bridging black-box learning and formal guarantees within a unified neurosymbolic framework.
1. Introduction
@@
Discrete-time dynamical systems offer a foundational framework for modeling and controlling a broad spectrum of digitally controlled autonomous systems, including autonomous vehicles, implantable medical devices, biological networks, and industrial process management. As these systems constitute a critical component of our safety infrastructure, high-assurance formal reasoning is essential to ensure their safe operation. A principled and efficient approach to providing such safety guarantees involves synthesizing barrier certificates [24, 1]—a functional analog of inductive invariants—whose existence serves as a formal proof of safety [23]. Although finding such certificates is inherently computationally challenging, recent neuro-symbolic approaches have shown considerable promise through the use of neural barrier certificates [7, 35, 34, 27]. While these neural certificates provide substantial flexibility in searching for safety guarantees, we posit that scalability and sample efficiency can be further improved in specialized control applications by identifying and leveraging structural properties of the underlying system. In this work, we focus on a subclass of discrete-time dynamical systems known as monotone systems [5, 29], which frequently arise in domains such as queuing networks [15], population dynamics [18], biological systems [4], and chemical networks [30]. By leveraging their structural characteristics, we synthesize barrier certificates represented as monotone neural networks, which we refer to as monotone neural barrier certificates.
@@
Monotone Systems. Monotone dynamical systems [29, 5] are characterized by the property that their trajectories preserve a natural partial order induced on their state set. Examples of monotone systems arise in domains such as biochemical reactions in synthetic biology [30], frequency regulation in power grids, congestion control in transportation networks [10, 8], population dynamics in ecology, and compartmental models in epidemiology [5]. Related concepts have also been studied in other domains, including positive linear systems [15], well-structured transition systems [16], and well-behaved transition systems [6]. It is well known that monotone systems exhibit highly ordered transient and asymptotic behaviors [29], making them particularly well-suited for data-driven methods, as their inherent structure helps mitigate the challenges posed by limited or noisy data. While monotonicity has informed control and stability analysis [26, 9], its potential for learning-based safety verification remains largely untapped.
@@
Barrier Certificates (A Proof of Safety). In safety verification, the central object of interest is the reachable set, which comprises all states the system can reach from a given set of initial states under a fixed control policy. Safety can be certified by showing that this set does not intersect the set of harmful or unsafe states. However, computing the exact reachable set is generally intractable for nonlinear or high-dimensional systems. A common alternative is to construct an inductive state-invariant, i.e., a set that contains the initial states and is closed under the system’s transitions. Barrier certificates [24] offer one such approach by defining a real-valued function over the state space whose level sets form a “barrier” separating safe and unsafe regions. Specifically, the barrier function must: 1) take values less than or equal to a constant over the initial states, 2) and strictly greater than that constant over unsafe states; and 3) decrease along system trajectories outside the unsafe set. A control barrier certificate (CBC) generalizes the idea of barrier certificate by requiring the existence of an action satisfying the conditions for a barrier certificate.
@@
Despite their intuitive appeal and ease of verification, complexity of finding CBCs remains a significant hurdle in their widespread adoption. Traditional methods for synthesizing CBCs rely on optimization-based formulations such as Sum-of-squares (SOS) programming [21, 17]. SOS is a popular approach that is applicable to systems with polynomial dynamics; however, it quickly becomes intractable due to the growth of semidefinite constraints with system dimension and polynomial degree. Linear programming (LP) methods, while more tractable, are typically limited to linear or affine certificates and cannot capture the rich nonlinear boundaries present in many real-world systems [28, 32]. Recent progress in deep learning has enabled the use of neural networks to be trained as barrier certificates [3, 19, 12, 20, 35, 31, 34]. These so-called neural control barrier certificates (NCBCs) represent both the barrier function and controller as neural networks. While expressive, these approaches often require full knowledge of the system dynamics (to verify correctness using constraint satisfaction tools such as Z3 [13]) or rely on bounding Lipschitz constants (which entails exponential sample complexity), limiting scalability and applicability to black-box systems [3, 34]. Moreover, these approaches treat the system as unstructured, ignoring properties like monotonicity that could drastically reduce the number of required samples and constraints.
@@
Monotone Neural Control-Barrier Certificates. We introduce an efficient scheme for synthesizing barrier certificates specifically tailored to (unknown) monotone systems. By exploiting monotonicity, we restrict the satisfaction of safety conditions to critical boundary slices of the state space—namely, the upper boundary of the initial set and the lower boundary of the unsafe set. This reduces the verification burden from the entire state space to a carefully selected subset—without compromising formal guarantees—and enables the use of gradient-based optimization to learn barrier certificates using simulation data alone. Our method avoids the need for explicit models, Lipschitz bounds, or semidefinite relaxations. This makes it uniquely suited for safety verification in large-scale, nonlinear systems where monotonicity arises naturally from physical or structural properties.
2. Preliminaries and Mathematical Foundations
@@
We denote the set of real, non-negative real, and positive real numbers by $\mathbb{R}$}, ${\mathbb{R}}_{\geq 0}$, and ${\mathbb{R}}_{>0}$, respectively. For sets $A$ and $B$, we use $A\setminus B$ for set difference and $A\times B$ for Cartesian product, and write $|A|$ for the cardinality of $A$. Given a vector $v=(v_{1},\ldots,v_{n})^{\top}\in{\mathbb{R}}^{n}$, its Euclidean norm is $\|v\|=\sqrt{\sum_{i=1}^{n}v_{i}^{2}}$. The rectified linear unit is defined by $\mathsf{ReLU}(x):=\max(x,0)$, and for vectors $x,y\in{\mathbb{R}}^{n}$, the mean squared error is $\mathsf{MSE}(x,y):=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-y_{i})^{2}$. We denote the $i$th component of a vector $x$ by $x_{i}$, and write $x\leq y$ to indicate element-wise inequality. A set $\mathcal{S}\subseteq{\mathbb{R}}^{n}$ is lower closed if $x\in\mathcal{S}$ and $y\leq x$ imply $y\in\mathcal{S}$; it is upper closed if $x\in\mathcal{S}$ and $y\geq x$ imply $y\in\mathcal{S}$. A hyper-rectangle is denoted by $\underline{x},\overline{x}]:=\{x\in{\mathbb{R}}^{n}\mid\underline{x}\leq x\leq\overline{x}\}$. Given a set $A\subseteq\mathbb{R}^{n}$, $x\in\mathbb{R}^{n}$ is a maximal (resp. minimal) point of $A$, if there does not exists any $y\in A$ such that $x\leq y$ (resp. $y\leq x$).
2.1. Safety of Dynamical Systems
@@
Definition 2.1 (Discrete-time dynamical system).
@@
(2.1) $x_{t+1}=f(x_{t},u_{t}),\qquad\text{for all }t\geq 0.$
The sequence $\{x_{t}\}_{t=0}^{\infty}$ defined by (2.1) is called a trajectory of $\mathfrak{S}$ for the input $\{u_{t}\}_{t=0}^{\infty}$. Given a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$, a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$, and an input sequence $\{u_{t}\}_{t=0}^{\infty}$, we say that $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$ for this input sequence if all trajectories starting from initial states in $\mathcal{X}_{0}$ remain outside of $\mathcal{X}_{u}$. That is, for every $x_{0}\in\mathcal{X}_{0}$, the corresponding trajectory $\{x_{t}\}_{t=0}^{\infty}$ for the input sequence $\{u_{t}\}_{t=0}^{\infty}$ starting from $x(0)=x_{0}$, satisfies
@@
In safety verification, inputs are viewed as disturbances, and the goal is to guarantee system safety under all possible input sequences $\{u_{t}\}_{t=0}^{\infty}\in U$. To this end, robust barrier certificates have been widely employed as an effective tool for safety verification [25].
Definition 2.2 (Robust barrier certificates).
@@
(2.2) $B(x)\leq\gamma,\qquad\text{ for all }x\in\mathcal{X}_{0},$
(2.3) $B(x)>\eta,\qquad\text{ for all }x\in\mathcal{X}_{u},\text{ and }$
(2.4) $B(x)\leq\delta\implies\mathrm{sup}_{u\in U}\{B(f(x,u))\}\leq\delta.$
The following theorem formulates the safety verification problem as a search for RBCs [25].
Theorem 2.3 (Safety verification via robust barrier certificates).
Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ and a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Suppose that there exists a robust barrier certificate satisfying conditions (2.2)-(2.4). For any input sequence $\{u_{t}\}_{t=0}^{\infty}\in U$, the system $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$.
@@
In safe control synthesis, inputs are treated as control policies, aiming to synthesize a controller $\{u_{t}\}_{t=0}^{\infty}$ that ensures system safety. Control barrier certificates have been widely adopted as an effective tool for synthesizing such controllers [2].
Definition 2.4 (Control barrier certificates).
Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. A function $B:\mathcal{X}\rightarrow{\mathbb{R}}$ is called a Control Barrier Certificate (CBC) for $\mathfrak{S}$ with respect to the unsafe set $\mathcal{X}_{u}\subseteq\mathcal{X}$ if there exists $\gamma,\delta,\eta\in{\mathbb{R}}_{\geq 0}$ such that $\gamma\leq\delta\leq\eta$ and $B$ satisfy conditions (2.2), (2.3), and the following condition
(2.5) $B(x)\leq\delta\implies\mathrm{inf}_{u\in U}\{B(f(x,u))\}\leq\delta.$
The next result formulates the safe control synthesis problem as a search for CBCs [2].
Theorem 2.5 (Control synthesis via control barrier certificates).
Consider a dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ and a set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Suppose that there exists a control barrier certificate satisfying conditions (2.2), (2.3), and (2.5). Then, there exists a map $K:\mathcal{X}\to U$ such that the system $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
2.2. Monotone Dynamical Systems